# Interpretable Visualizations of Data Spaces for Classification Problems

### v3

@@ I Introduction
Over the past decades, machine learning (ML) has transformed computational chemistry, for example accelerating materials discovery [1], improving the quality of molecular simulations [2, 3], and enabling reliable predictions of materials properties [4]. However, many traditional ML and deep learning methods function as “black boxes,” providing little insight into the underlying physics and chemistry of materials and materials processes. However, in chemical research, the true value of an ML model often goes beyond its ability to make accurate predictions; it also lies in its ability to reveal how the set of input parameters, often carefully chosen to capture the underlying chemistry, relate to the property of interest. Addressing this need for interpretability has driven the development of approaches specifically tailored to uncover structure-property relationships, a pursuit central to the field of Quantitative Structure-Property Relationship (QSPR) modeling.
Many approach QSPR through human-interpretable visualization via unsupervised machine-learning methods. Principal Component Analysis (PCA) is the most standard method for mapping tasks [5]. PCA involves identifying a linear (or non-linear, in the case of kernel PCA) combination of the features that maximizes the variance (i.e., information) retained. Beyond PCA, there exist a slue of other mapping techniques commonly used to qualitatively understand data spaces – t-distributed Stochastic Neighbor Embedding (t-SNE) [6], Uniform Manifold Approximation and Projection (UMAP) [7], and Isometric Mapping (ISOMAP) [8] focus on the preservation of the local, global, and geometric data structure, respectively, and Locally Linear Embedding (LLE) [9] and multi-dimensional scaling (MDS) [10, 11] preserve distances between neighboring data points. While all of these methods excel at creating informative visualizations and identifying patterns in the structure of the data, the “P” (property) of “QSPR” is only spuriously accounted for, where any relationship between the visualized dimensions and the property of interest is purely coincidental.
@@
Thus, while methods from unsupervised learning and supervised learning address aspects of QSPR, none can fully do so on their own. There has been an emergence of "visualization as interpretation" techniques that aim to provide a solution. Briefly, the vast majority of these methods aim to compute invertible mappings that represent the partitioning and clustering of data in higher-dimensional spaces [15, 16, 17, 18]. For example, DeepView, the most notable of these methods, combines UMAP and Fisher distance for projection and then samples the resultant latent space on a grid to assess data boundaries[19]. However, computational cost and model complexity have remained large challenges for many of these methods, particularly those reliant on multi-layer architectures or grid-based assessments. And, methods based on high-non-linear operations or architectures can result in topologically-bound latent spaces, where arbitrary points or pixels in a grid may not directly correspond to physically realizable data points [20].
We address these challenges by exploiting a family of methods, known as hybrid learning, that draws elements from both supervised and unsupervised learning. Some of these methods include principal components regression [21], partial least squares regression [22], cluster-wise regression [23], continuum regression [24], and principal covariates regression (PCovR) [25]. Of these methods, PCovR is particularly valuable for QSPR because it is directly formulated as a multi-objective problem in which the regression and dimensionality reduction tasks are optimized simultaneously. PCovR has broad applications within the fields of climate science [26], economic forecasting [27], psychology [28], bioinformatics [29, 30, 31], and molecular crystallization [32]. Similar to PCA, PCovR can also be used to incorporate non-linear relationships through similarity kernels [33], and can be used with regularization to create sparse solutions [26, 30]. In addition, PCovR has been used to modify other techniques, such as clusterwise regression [34], farthest point sampling, and CUR decomposition [35].
@@ II Theory
We provide a concise description of the underlying mathematics for principal covariates classification (PCovC). We direct interested readers to [33] for a more thorough discussion, including that of data scaling and incorporating non-linearity. Included in App. B are discussions of other QSPR techniques, namely SHAP analysis [13], linear discriminant analysis [36], and saliency maps [37].
@@ II.1 Notation
We assume that the data is organized into two objects: a “feature matrix”, $\mathbf{X}$, that contains a set of input variables for a set of observations, and a “target matrix”, $\mathbf{Y}$, that contains the classes for each observation. The rows of the feature matrix $\mathbf{X}$ correspond to the individual samples and the columns correspond to the features, leading to shape $n_{\text{samples}}\times n_{\text{features}}$. The target matrix $\mathbf{Y}$ has shape $n_{\text{{samples}}}\times n_{\text{{labels}}}$ and contains the output labels for each observation, each of which can take an integer value between 0 and $n_{\text{classes}}-1$. We denote the latent-space projection of the input data as $\mathbf{T}$, which has shape $n_{\text{samples}}\times n_{\text{components}}$. Finally, we denote $\mathbf{Z}$ as the evidence tensor, which contains confidence scores for each class. $\mathbf{Z}$ has shape of $n_{\text{samples}}\times n_{\text{classes}}\times n_{\text{labels}}$. We assume that $\mathbf{P_{XZ}}$ is determined by a trained classification model, where, depending on the architecture of this model, $\mathbf{Z}$ is used to determine the class of the data points. In this context, we wish to learn linear mappings between $\mathbf{X}$, $\mathbf{T}$, and $\mathbf{Z}$, which we can define using projection matrices. In this work, we define $\mathbf{P}_{AB}$ as the projector from space $\mathbf{A}$ to space $\mathbf{B}$. Therefore, we can write $\mathbf{T}\equiv\mathbf{XP}_{XT}$ and $\mathbf{Z}=\mathbf{TP}_{TZ}=\mathbf{XP}_{XT}\mathbf{P}_{TZ}.$ The relationship between these objects is shown in Fig. 1.
@@ II.2 Linear Classification
The goal of a classification model is to assign an unseen sample to one of $n_{\text{classes}}$ discrete classes. Thus, we aim to draw “decision boundaries” that divide the input space into $n_{\text{classes}}$ regions, with each region corresponding to a certain class. In a linear classification problem, these boundaries are lines or hyperplanes in a higher-dimensional space. Deterministic approaches, such as support vector machines, involve finding a discriminant function and use an activation function to map this function to a specific class label. Probabilistic approaches, such as logistic regression, estimate the probability distributions of the input and output and use these probabilities to make classifications. More generally, linear classifiers make predictions by establishing a linear projection of the input data $\mathbf{Z}=\mathbf{XP}_{XZ}.$ and then using some function $\phi$, typically called an activation function, to map the continuous $\mathbf{Z}$ to the set of discrete class labels $\mathbf{Y}\in{0,1,2,...n_{\text{classes}}-1}$, such that $\mathbf{Y}=\phi(\mathbf{Z}).$
@@ II.4 Principal Covariates Classification
The PCovR loss function, Eq. (1), should not be directly modified to be used in classification tasks – replacing the regression loss with a classification loss often results in a non-analytical function (when the underlying classification algorithm incorporates activation functions into their loss optimization) or is scientifically non-productive, as many classification tasks, particularly multiclass tasks, use numerical values to denote class, where these numerical values are ostensibly arbitrary. We choose to adapt PCovR for classification tasks by leveraging the evidence as our approximation of $\mathbf{Y}$. This yields a new modified Gram matrix (or modified covariance matrix)
@@
whose eigendecomposition can be used to obtain a relevant low-dimensional map. Because different classification algorithms operate to optimize different loss terms, they will yield different values of $\mathbf{Z}$; this is visualized in Fig. 2, where we demonstrate the new latent space maps determined for PCovC at $\alpha=0.5$ for the archetypal Iris dataset [38, 39] and several popular linear classification algorithms.
Similar to PCovR, PCovC will inherit the regularization incorporated into its parent supervised algorithm, which will prevent overfitting, and can be trivially adapted for non-linear kernel classification models, where instead $\mathbf{Z}\equiv\mathbf{KP}_{KZ}$. As written, this approach is not well-suited for decision trees or ensemble classification methods, which require multiple analogs of the evidence matrix.
@@ III.1 What makes a molecule toxic? Exposing decision boundaries in toxicity classifications
First, we compare the mapping results of PCovC as a function of our mixing parameter, $\alpha$, in Fig. 3. As expected, a PCA map of this dataset reflects little distinction between toxic and non-toxic molecules, where a subsequent analysis via logistic regression labels all data points as “non-toxic,” i.e., the majority population of the dataset. As shown in the SI, using nonlinear unsupervised dimensionality reduction algorithms such as t-SNE or UMAP also fails to distinguish toxic molecules from non-toxic molecules. As $\alpha$ is lowered and the classification task bears more weight in Eq. (1), the activity cliff becomes clearer and our predictive power improves, even superceding that of logistic regression on the full-dimensional data. It is worth noting that the best performance, unlike PCovR, comes from a lower $\alpha$ value than 0.5 – this is likely due to the class imbalance within this dataset, and demonstrates the benefits of the tunability of PCovC.
@@
Figure 4: Molecules from the training set near the decision boundary that demonstrate delineating features of toxicity based on the “tox21-ache-p3 (-1)” assay. The underlying map corresponds to PCovC at $\alpha=0.05$; markers denote different pairs of molecules, with blue and white markers corresponding to toxic and non-toxic molecules. Toxic/non-toxic molecule pairs were chosen to correspond to the molecules closest in PCovC space.
@@
Figure 5: Maps of the Tox21 dataset considering multiple assays for classification, two double-assay maps (left, center) and a triple-assay map (right). Each plot is given for the $\alpha$ that corresponds to the best classification accuracy for the corresponding assays. Marker color corresponds to the values for the three assays, shown in the table to the right. Opacity corresponds to the train/test split.
@@
Of the three assays, Assay 1 is the most sensitive indicator of toxicity. In our postprocessed dataset, 488 of the 5,767 (8.46%) compounds are labeled as toxic by this assay, compared to 193 (3.35%) and 146 (2.53%) compounds for the Assays 2 and 3, respectively. This means that there are significantly many more data points considered “toxic” by Assay 1 that are “non-toxic” according to Assays 2 and 3. The PCovC maps for these singular assays and their combinations are shown in Fig. 5. Note the number of red points, indicating molecules only considered “toxic” by Assay 1, and black points, indicating molecules considered toxic by all assays. A map generated from solely Assay 2 positions these red points with the “non-toxic” white points (as shown in the Supplementary Information). The clearest separation occurs in the map built on Assays 1 and 2, which delineates both decision boundaries clearly and discretely. One can imagine doing similar analysis to that of Fig. 4 to identify the molecules along these boundaries to determine the differentiation by the two assays.
@@ III.2 Parsing multi-class datasets: spectral analysis of organosulfur bonding environments
Figure 6: Analysis of organosulfur dataset for different classification targets. Rows correspond to classification based on oxidation state (top), bond type (center), and aromaticity (bottom). (left) Underlying XANES spectra, colored by the class delineations. The main figure corresponds to the underlying spectra. (right) Maps according to PCA, LDA, and PCovC, colored by the class delineations. The PCA map will be the same for all three analyses. In the upper right, we indicate the accuracy of a logistic regression model on the resulting 2D map for the offset testing data. For PCovC, $\alpha$ is chosen to optimize test classification accuracy, indicated in each figure.
@@
As can be seen in the first row of Fig. 6, it is straightforward to identify the the signals corresponding to different oxidation states, as the data separates well based on this set of classes. Because of this, a logistic regression model trained on the 2D PCA latent space yields perfect test accuracy, and LDA only misclassifies one test compound. Because PCA is well-suited to separate oxidation states, the choice of $\alpha$ for a PCovC map is somewhat arbitrary, hence we choose $\alpha=0.5$. Note that the map shown is qualitatively similar to the PCA projection, with some minute differences.
@@
Bond types are not well-separated in the PCA projection, given that the -2 oxidation state contains three different types. As shown in Fig. 6, the sulfides and the thiols are the most difficult to separate, leading to a loss in accuracy for both PCA (87%) and LDA (85%) . However, a logistic regression model trained on the PCovC latent space is still able to achieve a 92% bond type test classification accuracy. This demonstrates how incorporating classification loss into our mapping procedures “refocuses” our model; in purely unsupervised schema, having a wealth of data within one class versus another leads to this data dominating the mapping task.
@@
The most difficult task is separating the compounds by aromaticity, as this is a property that cross-cuts the “higher-order” classes. As is suggested by the maps in Fig. 6, the principal components seem to be capturing the oxidation state and bond type of the organosulfur compounds. However, because neither of these properties has a direct relationship with bond aromaticity, a logistic regression model trained from the PCA latent space struggles (accuracy of 48%). However, PCovC and LDA are each able to find latent spaces in which the compounds are well delineated by their aromaticity, with accuracies of 77% and 73%, respectively.
@@
Figure 7: Projections of inorganic materials found in [46] using Kernel PCA, KDA, and Kernel PCovC, and the corresponding $\alpha$ analysis for this non-linear mapping. Color and opacity denote metal (blue) versus non-metal (red) and training (transparent) versus testing (opaque) splits, respectively. Above each map is the distribution of points in each class.
@@
III.3 Understanding feature importance in non-linear classification of metals/nonmetals
Finally, what if each of our features are physically meaningful but we require non-linear models for effective classification? In this instance, it is unclear how these meaningful features contribute to the learned weights of a classification algorithm. However, with robust structure-property maps, we can still interpret these non-linear feature relationships. In this case study, we apply a Kernelized PCovC (KPCovC) on the dataset of 4,916 inorganic materials published in [46] to understand the relationships between elemental properties and metallic behavior. Each of these materials is represented as a set of 136 properties of the constituent elements, and the learnable target is whether the material is a metal or a non-metal. Following the procedure in [46], we define a non-metal and metal as materials that do or do not have an electronic band gap, and extract the evidence matrix from a trained support vector classification (SVC) model, where $\mathbf{Z}\equiv\mathbf{KP}_{KZ},$ where $\mathbf{K}$ is a similarity matrix determined by a radial basis function (RBF) kernel, using a Gaussian width $\gamma$ of 0.01.
With this kernel matrix, we compute the Kernel PCA, Kernel Discriminant Analysis (KDA) and KPCovC projections of the elemental property data, shown in Fig. 7. As was shown in [46], we also observe that applying KPCovC with a support vector machine classifier optimizes classification performance and best exposes the decision boundary; using KPCovC with $\alpha=0.1$, the accuracy to 90% compared to 88% and 65% with the analogous KDA and KPCA, respectively.
Figure 8: Correlations of materials characteristics with the non-linear KPCovC and KPCA maps. (left) Absolute value of Pearson correlation coefficients for different input features with the resulting maps from Fig. 7. Darker squares denote stronger correlation. (middle) Ranked SHAP values and random forest feature importance. Darker squares indicate a higher importance. (right) Recoloring of KPCovC (main) and KPCA (inset) by feature values, from low (red) to high (blue) on a linear scale.

### v4

@@ I Introduction
Over the past decades, machine learning (ML) has transformed computational chemistry, for example accelerating materials discovery [30], improving the quality of molecular simulations [4, 5], and enabling reliable predictions of materials properties [37]. However, many traditional ML and deep learning methods function as “black boxes,” providing little insight into the underlying physics and chemistry of materials and materials processes. In chemical research, the true value of an ML model often goes beyond its ability to make accurate predictions; it also lies in its ability to reveal how the set of input parameters, often carefully chosen to capture the underlying chemistry, relate to the property of interest. Addressing this need for interpretability has driven the development of approaches specifically tailored to uncover structure-property relationships, a pursuit central to the field of Quantitative Structure-Property Relationship (QSPR) modeling.
Many approach QSPR through human-interpretable visualization via unsupervised machine-learning methods. Principal Component Analysis (PCA) is the most standard method for mapping tasks [35]. PCA involves identifying a linear (or nonlinear, in the case of kernel PCA) combination of the features that maximizes the variance (i.e., information) retained. Beyond PCA, there exist a slue of other mapping techniques commonly used to qualitatively understand data spaces – t-distributed Stochastic Neighbor Embedding (t-SNE) [32, 38], Uniform Manifold Approximation and Projection (UMAP) [33], and Isometric Mapping (ISOMAP) [47] focus on the preservation of the local, global, and geometric data structure, respectively, and Locally Linear Embedding (LLE) [41] and multi-dimensional scaling (MDS) [25, 26] preserve distances between neighboring data points. While all of these methods excel at creating informative visualizations and identifying patterns in the structure of the data, the “P” (property) of “QSPR” is only spuriously accounted for, where any relationship between the visualized dimensions and the property of interest is purely coincidental.
@@
Thus, while methods from unsupervised learning and supervised learning address aspects of QSPR, none can fully do so on their own. Consequently, this problem has motivated the extension of existing unsupervised learning algorithms to incorporate target information into their mapping strategies, with two beneficiaries of this direction having been UMAP and PCA. With UMAP, supervision can be added by intersecting the unsupervised UMAP graph constructed on the input features with a graph constructed on the targets. However, recent works have validated that because unsupervised UMAP is already designed for resolving clusters, adding supervision does not significantly alter performance [58]. Supervised PCA, on the other hand, involves modifying the principal components to instead identify directions in which the dependence between the input and output data is maximized [3, 16]. More specifically, supervised PCA learns a linear (or kernel, in kernel supervised PCA) projection that maximizes an empirical approximation of the Hilbert-Schmidt independence criterion [18], which itself is a function of the cross-covariance between the input and output data. While this formulation ensures that classes would be well-separated, its major shortcoming is that it does not explicitly preserve the variance of the input data. This can result in cases in which directions in the data that are correlated with the target due to collinearity are overrepresented in the map, despite the fact that these directions are likely noise.
Concurrently, there has been an emergence of “visualization as interpretation” techniques that aim to provide a solution. Briefly, the vast majority of these methods aim to compute invertible mappings that represent the partitioning and clustering of data in higher-dimensional spaces [52, 40, 34, 6]. For example, DeepView, the most notable of these methods, combines UMAP and Fisher distance for projection and then samples the resultant latent space on a grid to assess data boundaries [42]. However, computational cost and model complexity have remained large challenges for many of these methods, particularly those reliant on multi-layer architectures or grid-based assessments. And, methods based on high-nonlinear operations or architectures can result in topologically-bound latent spaces, where arbitrary points or pixels in a grid may not directly correspond to physically realizable data points [53].
We address these challenges by exploiting a family of methods, known as hybrid learning, that draws elements from both supervised and unsupervised learning. Some of these methods include principal components regression [24], partial least squares regression [55], cluster-wise regression [44], continuum regression [45], and principal covariates regression (PCovR) [12]. Of these methods, PCovR is particularly valuable for QSPR because it is directly formulated as a multi-objective problem in which the regression and dimensionality reduction tasks are optimized simultaneously. PCovR has broad applications within the fields of climate science [13], economic forecasting [20], psychology [50], bioinformatics [10, 49, 46], and molecular crystallization [9]. Similar to PCA, PCovR can also be used to incorporate nonlinear relationships through similarity kernels [21], and can be used with regularization to create sparse solutions [13, 49]. In addition, PCovR has been used to modify other techniques, such as clusterwise regression [54], farthest point sampling, and CUR decomposition [8].
@@ II Theory
We provide a concise description of the underlying mathematics for principal covariates classification (PCovC). We direct interested readers to [21] for a more thorough discussion, including that of data scaling and incorporating nonlinearity. Included in App. B are discussions of other QSPR techniques, namely SHAP analysis [31], linear discriminant analysis [19], and saliency maps [43].
@@ II.1 Notation
We assume that the data is organized into two objects: a “feature matrix”, $\mathbf{X}$, that contains a set of input variables for a set of observations, and a “target matrix”, $\mathbf{Y}$, that contains the classes for each observation. The rows of the feature matrix $\mathbf{X}$ correspond to the individual samples and the columns correspond to the features, leading to shape $n_{\text{samples}}\times n_{\text{features}}$. The target matrix $\mathbf{Y}$ has shape $n_{\text{{samples}}}\times n_{\text{{labels}}}$ and contains the output labels for each observation, each of which can take an integer value between 0 and $n_{\text{classes}}-1$. We denote the latent-space projection of the input data as $\mathbf{T}$, which has shape $n_{\text{samples}}\times n_{\text{components}}$. Finally, we denote $\mathbf{Z}$ as the evidence tensor, which contains confidence scores for each class. $\mathbf{Z}$ has shape of $n_{\text{samples}}\times n_{\text{classes}}\times n_{\text{labels}}$. We assume that $\mathbf{P_{XZ}}$ is the set of weights learned by a trained classification model, where, depending on the architecture of this model, $\mathbf{Z}$ is used to determine the class of the data points. In this context, we wish to learn linear mappings between $\mathbf{X}$, $\mathbf{T}$, and $\mathbf{Z}$, which we can define using projection matrices. In this work, we define $\mathbf{P}_{AB}$ as the projector from space $\mathbf{A}$ to space $\mathbf{B}$. Therefore, we can write $\mathbf{T}\equiv\mathbf{XP}_{XT}$ and $\mathbf{Z}=\mathbf{TP}_{TZ}=\mathbf{XP}_{XT}\mathbf{P}_{TZ}.$ The relationship between these objects is shown in Fig. 1.
@@ II.2 Linear Classification
The goal of a classification model is to assign an unseen sample to one of $n_{\text{classes}}$ discrete classes. Thus, we aim to draw “decision boundaries” that divide the input space into $n_{\text{classes}}$ regions, with each region corresponding to a certain class. In a linear classification problem, these boundaries are lines or hyperplanes in a higher-dimensional space. Deterministic approaches, such as support vector machines, involve finding a discriminant function and use an activation function to map this function to a specific class label. Probabilistic approaches, such as logistic regression, estimate the probability distributions of the input and output and use these probabilities to make classifications. More generally, linear classifiers make predictions by learning a linear projection of the input data $\mathbf{Z}=\mathbf{XP}_{XZ},$ and then using some function $\phi$, typically called an activation function, to map the continuous $\mathbf{Z}$ to the set of discrete class labels $\mathbf{Y}\in{0,1,2,...n_{\text{classes}}-1}$, such that $\mathbf{Y}=\phi(\mathbf{Z}).$
@@ II.4 Principal Covariates Classification
The PCovR loss function, Eq. (1), should not be directly modified to be used in classification tasks – replacing the regression loss with a classification loss often results in a non-analytical function (when the underlying classification algorithm incorporates activation functions into their loss optimization) or is scientifically non-productive, as many classification tasks, particularly multiclass tasks, use numerical values to denote class, where these numerical values are ostensibly arbitrary. We choose to adapt PCovR for classification tasks by leveraging the evidence as our approximation of $\mathbf{Y}$. This yields a new objective function
[…]
and a new modified Gram matrix (or modified covariance matrix)
@@
whose eigendecomposition can be used to obtain a relevant low-dimensional map. Similar to PCovR, this procedure can be adapted for nonlinear kernel classification models, where instead $\mathbf{Z}\equiv\mathbf{KP}_{KZ}$.
Thus, an important component of this procedure is the selection of the expectation matrix $\mathbf{Z}$. As different classification algorithms operate to optimize different loss terms, they will yield different values of $\mathbf{Z}$; this is visualized in Fig. 2, where we demonstrate the new latent space maps determined for PCovC at $\alpha=0.5$ for the archetypal Iris dataset [14, 1] and several popular linear classification algorithms. These maps are well-founded, as, by construction, there exists a strong relationship between $\mathbf{X}$ and $\mathbf{Z}$. This framework indeed is strongest where there exists a clear relationship between $\mathbf{X}$ and $\mathbf{Z}$ (or, equivalently, $\mathbf{K}$ and $\mathbf{Z}$). Mathematically, this allows the easy identification of covariances between variables, and thus the loss written in Eq. 1 to vary smoothly with $\alpha$.
However, this does not impose a necessary requirement that $\mathbf{Z}$ comes from a classification model constructed by $\mathbf{Y}=\phi(\mathbf{Z}),$ simply that $\mathbf{Z}$ reflects the class likelihoods and is normalized such that its variance is of similar magnitude to $\mathbf{X}$ (having $\mathbf{X}$ and $\mathbf{Z}$ of non-standardized variance leads to asymptotic behavior of relevant loss functions with respect to $\alpha$, as discussed in Helfrecht et al. [21]). Also, $\mathbf{Z}$ should be determined in such a way as to prevent over-fitting, as is typically done by regularizing losses in classification models.
Choosing $\alpha$ depends on the intention behind the mapping procedure (how much one cares about the reconstruction of $\mathbf{X}$ versus the clear visualization of the decision boundary), but can be done programmatically based on the mixed loss (Eq. 2). When the intention is to understand the behavior of the classifier, it is important to note that for binary classification with $\mathbf{Z}=\mathbf{XP}_{XZ}$ or $\mathbf{Z}=\mathbf{KP}_{KZ}$, $\mathbf{Z}$ should be fully reproducible from $\mathbf{T}$ as $\alpha\to 0$. This makes low (but nonzero) $\alpha$ values, such as $0.05$, more attractive, as the classifier’s learned decision boundary can be well-captured in this regime while also learning an orthogonal dimension for analysis. With multiclass problems, we require that $\mathbf{T}$ has at least $n_{\text{classes}}$ dimensions to fully reconstruct $\mathbf{Z}$, thus the same recommendations are not directly applicable.
With perfectly balanced datasets wherein the $\mathbf{X}$ and $\mathbf{Z}$ matrices are standardized to unit variance, $\alpha$ can typically be taken as 0.5, as this weights the reconstruction of the matrices equally. When the matrices are of different variances or the classes are imbalanced, $\alpha$ should be determined by a minimization of PCovC loss, ideally through cross-validation. We have followed this procedure in all subsequent case studies, and have included corresponding cross-validation results in the Supplementary Information. For a more thorough discussion of the impact of $\alpha$, we direct the interested reader to Vervloet et al. [51].
@@ III.1 What makes a molecule toxic? Exposing decision boundaries in toxicity classifications
First, we compare the mapping results of PCovC as a function of our mixing parameter, $\alpha$, in Fig. 3. As expected, a PCA map of this dataset reflects little distinction between toxic and non-toxic molecules, where a subsequent analysis via logistic regression labels all data points as “non-toxic,” i.e., the majority population of the dataset. As shown in App. D, nonlinear unsupervised dimensionality reduction algorithms such as t-SNE or UMAP also fail to distinguish toxic molecules from non-toxic molecules. As $\alpha$ is lowered and the classification task bears more weight in Eq. (1), the activity cliff becomes clearer and our predictive power improves, even superceding that of logistic regression on the full-dimensional data.
@@
Figure 4: Molecules from the training set near the decision boundary that demonstrate delineating features of toxicity based on the “tox21-ache-p3 (-1)” assay. The underlying map corresponds to PCovC at $\alpha=0.05$; markers denote different pairs of molecules, with blue and white markers corresponding to toxic and non-toxic molecules. Toxic/non-toxic molecule pairs were chosen to correspond to the molecules closest in PCovC space. Atoms are colored using the standardized Corey-Pauling-Koltun coloring scheme.
@@
Figure 5: PCovC maps of the Tox21 dataset considering multiple assays for classification, two double-assay maps (left, center) and a triple-assay map (right). Marker color corresponds to the values for the three assays, shown in the table to the right. Opacity denotes train (translucent)/ test (opaque) split.
@@
Of the three assays, Assay 1 is the most sensitive indicator of toxicity. In our postprocessed dataset, 488 of the 5,767 (8.46%) compounds are labeled as toxic by this assay, compared to 193 (3.35%) and 146 (2.53%) compounds for the Assays 2 and 3, respectively. This means that there are significantly many more data points considered “toxic” by Assay 1 that are “non-toxic” according to Assays 2 and 3. The PCovC maps for these singular assays and the other assay combinations not shown in Fig. 5 are shown in the Supplementary Information. Note the number of red points, indicating molecules only considered “toxic” by Assay 1, and black points, indicating molecules considered toxic by all assays. A map generated from solely Assay 2 positions these red points with the “non-toxic” white points (as shown in the Supplementary Information). The clearest separation occurs in the map built on Assays 1 and 2, which delineates both decision boundaries clearly and discretely. One can imagine doing similar analysis to that of Fig. 4 to identify the molecules along these boundaries to determine the differentiation by the two assays.
@@ III.2 Parsing multi-class datasets: spectral analysis of organosulfur bonding environments
Figure 6: Analysis of organosulfur dataset for different classification targets. Rows correspond to classification based on oxidation state (top), bond type (center), and aromaticity (bottom). (left) Underlying XANES spectra, colored by the class delineations. The main figure corresponds to the underlying spectra. (right) Maps according to the first and second learned PCA, LDA, and PCovC dimensions, colored by the class delineations. The PCA map will be the same for all three analyses. In the upper right, we indicate the accuracy of a logistic regression model on the resulting 2D map for the offset testing data.
@@
As can be seen in the first row of Fig. 6, it is straightforward to identify the the signals corresponding to different oxidation states, as the data separates well based on this set of classes. Because of this, a logistic regression model trained on the 2D PCA latent space yields perfect test accuracy, and LDA only misclassifies one test compound. Because PCA is well-suited to separate oxidation states, the choice of $\alpha$ for a PCovC map is somewhat arbitrary, with $\alpha=0.19$ minimizing the PCovC loss. Note that the map shown is qualitatively similar to the PCA projection, with minute differences.
@@
Bond types are not well-separated in the PCA projection, given that the -2 oxidation state contains three different types. As shown in Fig. 6, the sulfides and the thiols are the most difficult to separate, leading to a loss in accuracy for both PCA (87%) and LDA (85%) . However, a logistic regression model trained on the PCovC latent space is still able to achieve a 89% bond type test classification accuracy, despite the fact that the $\alpha$ value was not chosen to maximize it. This demonstrates how incorporating classification loss into our mapping procedures “refocuses” our model; in purely unsupervised schema, having a wealth of data within one class versus another leads to this data dominating the mapping task.
@@
The most difficult task is separating the compounds by aromaticity, as this is a property that cross-cuts the “higher-order” classes. As is suggested by the maps in Fig. 6, the principal components seem to be capturing the oxidation state and bond type of the organosulfur compounds. However, because neither of these properties has a direct relationship with bond aromaticity, a logistic regression model trained from the PCA latent space struggles (accuracy of 48%). However, PCovC and LDA are each able to find latent spaces in which the compounds are well delineated by their aromaticity, with accuracies of 77% and 73%, respectively. The PCovC latent space also captures the difference in bond type via the second principal covariate, where the upper cluster of points correspond to the sulfones, and the lower cluster encompasses the sulfoxides, thiocarbonyls, sulfides, and thiols.
@@
Figure 7: Projections of inorganic materials found in [59] using Kernel PCA, KDA, and Kernel PCovC. Color and opacity denote metal (blue) versus non-metal (red) and training (transparent) versus testing (opaque) splits, respectively. Above each map is the distribution of points in each class.
@@
III.3 Understanding feature importance in nonlinear classification of metals/nonmetals
Finally, what if each of our features are physically meaningful but we require nonlinear models for effective classification? In this instance, it is unclear how these meaningful features contribute to the learned weights of a classification algorithm. However, with robust structure-property maps, we can still interpret these nonlinear feature relationships. In this case study, we apply a Kernelized PCovC (KPCovC) on the dataset of 4,916 inorganic materials published in [59] to understand the relationships between elemental properties and metallic behavior. Each of these materials is represented as a set of 136 properties of the constituent elements, and the learnable target is whether the material is a metal or a non-metal. Following the procedure in [59], we define a non-metal and metal as materials that do or do not have an electronic band gap, and extract the evidence matrix from a trained support vector classification (SVC) model, where $\mathbf{Z}\equiv\mathbf{KP}_{KZ},$ where $\mathbf{K}$ is a similarity matrix determined by a radial basis function (RBF) kernel, using a Gaussian width $\gamma$ of 0.01.
With this kernel matrix, we compute the Kernel PCA, Kernel Discriminant Analysis (KDA) and KPCovC projections of the elemental property data, shown in Fig. 7. As was shown in [59], we also observe that applying KPCovC with a support vector machine classifier optimizes classification performance and best exposes the decision boundary; using KPCovC with $\alpha=0.12$, the accuracy to 90% compared to 88% and 65% with the analogous KDA and KPCA, respectively.
Figure 8: Correlations of materials characteristics with the nonlinear KPCovC and KPCA maps. (left) Absolute value of Pearson correlation coefficients for different input features with the resulting maps from Fig. 7. Darker squares denote stronger correlation. (middle) Ranked SHAP values and random forest (RF) feature importance. Darker squares indicate a higher importance.
(right) Recoloring of KPCovC (main) and KPCA (inset) by feature values, from low (red) to high (blue) on a linear scale.